OpenStudy (anonymous):

Tracy Finely obtained a loan of $2,500 to buy a computer and a printer. The interest rate is 8.5% and the monthly payment over 24 months is $113.64. How much of the first monthly payment covered principal? I have to use the formula I=PRT I plugged in the numbers : I=$2,500x8.5%x24 But my answer isn't right. Even after I subtracted.

OpenStudy (wolf1728):

I calculate $17.71 is that right?

OpenStudy (anonymous):

How did you get $17.71? Once I subtract it from $113.64-$17.71=$95.93. Which I think is the answer.

OpenStudy (wolf1728):

That is the interest rate amount the principal is $95.93 right?? I guess it is!!!

OpenStudy (anonymous):

But each time I plugged in my number I never got the $17.71. I want to know how you got it before I go.

OpenStudy (wolf1728):

Okay for one thing I have a website (all free) that explains these things (and loans and mortgages are a real pain) I'll explain it

OpenStudy (anonymous):

Okay, cool thank you!

OpenStudy (wolf1728):

First of all were you required to calculate the monthly payment ?

OpenStudy (anonymous):

Ahhh!!! I didn't put the 2/24 before multiplying it out. Which is how you got your answer!

OpenStudy (anonymous):

I thought it was 1/24 or 1/12

OpenStudy (wolf1728):

Anyway, my month by month loan calculator is located here: http://www.1728.org/mortmnts.htm

OpenStudy (anonymous):

Thank you for your help and for the website I'm sure it will come into use with my homework!

OpenStudy (wolf1728):

Notice that the amount going to interest and the amount paying principal changes every MONTH.

OpenStudy (anonymous):

I'll make note of it!

OpenStudy (wolf1728):

Do you need to see how to find the first interest payment and principal?

OpenStudy (anonymous):

I don't think so, I think I've got it. If I need help again do you mind if I tag you?

OpenStudy (wolf1728):

You may go right ahead and do so. (As I said loans, mortgages, credit cards are VERY difficult aren't they?)

OpenStudy (anonymous):

Yess they are! Ugh haha!

OpenStudy (wolf1728):

Okay, well good luck with those problems. :-)

OpenStudy (anonymous):

Thank you :). Have a good morning!

OpenStudy (wolf1728):

You too

OpenStudy (wolf1728):

Anyway, to see what goes to interest and what goes to equity, here's a good explanation: http://www.1728.org/loanmath.htm Anyway we must convert 8.5 % to a MONTHLY payment .085 / 12 = 0.0070833333 Interest Paid = 2,500 * 0.0070833333 = 17.71 Equity = Monthly Payment -17.71 = 95.93 New Balance = Previous Balance - Equity New Balance = 2,500 -95.93 = 2,404.07